Circular No. 5741

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 E. Magnier and W. Lewin, Masachusetts Institute of Technology; L.

Lubin, Princeton University; J. van Paradijs, University of Amsterdam; U. Zimmerman, Max Planck Institute for Extraterrestrial Physics; and G. Fabbiano, Center for Astrophysics, report: “We have found an object within 0′′.25 of the position of SN 1993J CCD images taken 1992 Jan. 24 at the Michigan-Dartmouth-MIT Observatory’s 2.4-m Hiltner telescope. The object appears stellar (FWHM 1′′.8) but is in a somewhat crowded region. We find the following magnitudes and colors: V = 19.7±0.5, V −R = +0.8±0.1, m(Hα) = 19.54 ± 0.1, m([S II]) = 19.85 ±0.1. The V magnitude is a quicklook result, and the error should reduce to < 0.1 mag with careful calibration. The emission-line fluxes are from 3.5-nm-wide filter images centered at 655.0 and 673.0 nm, and magnitudes are calibrated using the standard stars of Oke and Gunn (1983, Ap.J. 266, 713). The colors do not appear to be those of an emission-line object.”

J. Salzer, W. Herbst and G. Vinton, Wesleyan University, report: “A CCD image taken by Salzer with the 0.9-m telescope at Kitt Peak on 1992 Oct. 2 includes the field of SN 1993J. A single R image of 300-s duration was obtained under photometric conditions and shows a faint, stellar-appearing object with R = 20.0 ± 0.2 at the supernova’s position. CCD images obtained with the Van Vleck Observatory’s 0.6-m reflector on 1993 Mar. 31.2 UT confirm the identity of the supernova progenitor to within 0′′.2. The nearby star GSC 4383.0340 has R = 14.14 ± 0.04, and the supernova was at R = 10.7 in the VVO observations. The observed progenitor brightness is 0.7 mag fainter than the Mar. 1990 measurement reported on IAUC 5736, suggesting significant light variations in the progenitor—as would be expected in a late-type supergiant. A distance modulus of 27.6 gives a luminosity for the progenitor of M(V) = −6.9, consistent with a K0 Ia.”

G. Calamai, G. Govazzi, I. Randone and G. Tofani report that nearinfrared photometry at the CNR Tirgo Observatory, Gornergrat, gave J = 10.57 ± 0.05, H = 10.48 ± 0.05, K = 10.41 ± 0.05, in less than perfect sky conditions during Mar. 31.83-31.93 UT. Photoelectric photometry by G. Sostero, A. Lepardo and U. Munari with the Friuli Astronomical-Meteorological Association’s 0.45-m telescope on Mar. 31.972 UT gave V = 11.06, B − V = −0.35 [sic] (±0.02, comparison HD 86677). CCD photometry by P. Prugniel at Haute Provence on Apr. 1.00 UT gave B = 11.40, B − R = +0.45 (±0.03, Cousins system; cf. IAUC 5736).

1993 April 2 Brian G. Marsden